# On the Fundamental Limits of LLMs at Scale

### v1

@@ 2.3 Evaluation Misalignment and Guessing Incentives
The vast majority of present-day LLM evaluations employ binary grading. Responses are scored as either correct (1) or incorrect (0), with no partial credit for expressing uncertainty Xu et al. (2024a). Under such schemes, abstentions, which can be responses like “I don’t know,” “I’m uncertain,” or “I cannot answer without more information”, receive the same zero score as confident but incorrect answers. This creates a rational incentive structure that strictly favors guessing over abstention.
@@
A meta-analysis of influential benchmarks confirms this pervasive issue Yang et al. (2023). Among the most widely cited evaluations, namely MMLU-Pro, GPQA, Omni-MATH, IFEval, SWE-bench, MATH, BBH, HLE, all employ binary grading with no credit for uncertainty expressions Li et al. (2023c). MMLU-Pro and GPQA are standard multiple-choice exams with no “I don’t know” option. Omni-MATH uses equivalence grading (often via LM judges) that compares outputs to ground-truth answers, assigning full credit for correctness and zero otherwise. IFEval grades instruction-following compliance programmatically, again binarized. Only WildBench, which evaluates real user chats on a 10-point scale, offers partial credit for uncertainty, but even there, the rubric suggests that “I don’t know” responses score lower than “fair” responses with minor hallucinations, thus still incentivizing guessing Kirichenko et al. (2025).
@@
When models are trained via reinforcement learning from human feedback (RLHF) or direct preference optimization (DPO), the reward model inherits biases from human annotators who often prefer confident, fluent answers over hedged or uncertain ones Xu et al. (2024a). Even when a model is unsure, producing a plausible-sounding fabrication receives higher reward than admitting ignorance. This creates a reward hacking dynamic, wherein models learn to maximize perceived confidence and fluency, which correlates with reward, rather than factual accuracy.
@@
where $\beta$ is the inverse temperature. If $R$ rewards verbosity and confidence over accuracy, the learned policy $\pi^{*}$ shifts toward hallucination-prone policies that generate detailed fabrications. This phenomenon is exacerbated by LM-as-judge evaluation, where another LLM grades outputs. LM judges are susceptible to length bias, favoring longer responses even when they contain errors, and often fail to detect subtle factual inaccuracies, grading incorrect but fluent hallucinations as correct Yao et al. (2024a).
@@
In practice, LLMs exhibit significant overconfidence. They assign high probabilities to incorrect answers, i.e., $\mathbb{P}[\text{correct}\mid p_{\theta}(c)=0.9]\ll 0.9$. This miscalibration is particularly severe on long-tail queries and out-of-distribution inputs, where the model has minimal training signal Sainz et al. (2023).
@@
where $\mathcal{R}_{\text{mem}}(h)\ll\mathcal{R}_{\text{true}}(h)$ reflects near-perfect recall of memorized answers. This distorts leaderboard rankings: models with higher contamination rates appear to hallucinate less, while their real-world performance remains poor Xu et al. (2024a).
@@
Addressing evaluation misalignment requires major socio-technical reforms Kirichenko et al. (2025). First, confidence-aware grading should assign partial credit for well-calibrated uncertainty expressions. Instead of binary $\{0,1\}$ scores, evaluations could use:
@@ 2.4 Creativity-Factuality Trade-off
LLMs generate text by sampling from a learned probability distribution $p_{\theta}(x_{t}\mid x_{<t})$ over next tokens. The sampling strategy governs the balance between exploitation (selecting high-probability, safe continuations) and exploration (venturing into lower-probability, generative territory). This trade-off is controlled by hyperparameters like temperature $T$ and nucleus (top-$p$) sampling Peeperkorn et al. (2024).
@@
where $z_{t}$ is the logit for token $x_{t}$. As $T\to 0$, sampling becomes deterministic (greedy decoding), favoring the single most likely token. As $T\to\infty$, the distribution flattens toward uniform (high exploration, minimal exploitation). Empirically, low temperatures ($T\approx 0.1$-$0.5$) produce repetitive, conservative text with high factual accuracy but limited novelty. High temperatures ($T\approx 1.0$-$2.0$) yield diverse, creative outputs but also frequent hallucinations as the model samples unlikely tokens that lead to fabricated details Nguyen et al. (2024).
@@
The challenge is distinguishing these cases. Models lack explicit mechanisms to estimate epistemic vs. aleatoric uncertainty, so sampling strategies cannot adapt. A fixed high temperature boosts creativity but also hallucination, while a fixed low temperature reduces hallucination but also stifles creativity Farquhar et al. (2024).
@@
showing a negative correlation: improving creativity ($d\mathcal{C}>0$) necessitates reducing accuracy ($d\mathcal{A}<0$), and vice versa Nguyen et al. (2024).
@@
A provocative theoretical perspective argues that hallucination is necessary for improvisation in LLMs Jiang et al. (2024). If a model were constrained to produce only outputs directly supported by its training data, it could never generate truly novel text because every sequence then would be a recombination of observed patterns. Creativity demands extending beyond the training distribution, which by definition involves assigning non-zero probability to unseen or low-probability continuations. These exploratory samples constitute hallucinations when they introduce factual errors, but they are indispensable for creative generation. Formally, let $\mathcal{D}_{\text{train}}$ denote the support of the training distribution. A model that generates only from $\mathcal{D}_{\text{train}}$ produces zero hallucinations but also zero novel outputs. To improvise, the model must sample from regions where $p_{\text{train}}(x)\approx 0$ but $p_{\theta}(x)>0$. Whether such samples are creative or hallucinatory depends on the task: in fiction, they are celebrated as originality; in fact-based QA, they are condemned as fabrications. The line between creativity and hallucination blurs, with both stemming from the model’s learned ability to extrapolate beyond its training data.
@@ 3.1 Problem Description
Transformer-based LLMs have pushed context lengths into the tens or even hundreds of thousands of tokens, enabling applications like book summarization, multi-document QA, and long conversation memory Vaswani et al. (2017); Beltagy et al. (2020a); Zaheer et al. (2020); OpenAI Achiam et al. (2023); Grattafiori et al. (2024); Liu et al. (2024a); Tomczak and Kuppannagari (2025). Yet despite larger context windows, practical long-context reasoning often falls short of expectations Qiu et al. (2020); Han et al. (2021). Models frequently cannot effectively utilize the full window. The effective usable context of many LLMs is much shorter than their nominal context length An et al. (2024); Liu et al. (2023c); An et al. ; Huang et al. (2023). Even a 70B model trained to 128K context (Llama3.1) was found to only leverage about 64K effectively Grattafiori et al. (2024); An et al. and in a comprehensive benchmark (LongBench) with average inputs  6.7K words, even a strong 16K-token model (GPT-3.5-Turbo-16k) still struggles on longer contexts Radford et al. (2019); Brown et al. (2020); Wei et al. (2022). Clearly, simply extending the context window does not guarantee strong reasoning across that entire length. This review explores the reason why. We focus on three core factors (as shown in Figure 5), including training data positional distribution, positional encoding limits, and attention computation constraints, and explain how each fundamentally hinders long-range reasoning An et al. ; Huang et al. (2023).
@@ 3.2 Left-Skewed Training Distribution and Undertraining of Long Positions
One root cause is the left-skewed distribution of token positions in training data An et al. . During pretraining, model inputs are far more likely to be shorter texts or early segments of long texts, rather than full-length sequences. This creates a severe imbalance: tokens at later positions (far right in the context) are extremely underrepresented Xiong et al. (2023), as also shown in Figure 4. Recent analyses confirm this “left-skewed position frequency distribution” in large corpora An et al. (2024; ). For example, in the SlimPajama dataset (a massive webtext corpus), the frequency of examples using very long-range token distances is vanishingly low: using a 2048 token window, less than 20% of the training pairs involve distances in the upper half of the window, and less than 5% involve the extreme end of the window Bai et al. (2024a).
@@
In gradient-based training, rarely seen position interactions receive only tiny updates An et al. ; Liu et al. (2023c). Intuitively, the model learns to predict the next token mainly using nearby context, and it gets much less practice using very distant context. Formally, one can consider the training loss as an expectation over position indices: if the probability of a position beyond, say, 75% of the max context is extremely small, then the contribution of those positions to the loss (and thus to gradient updates) is negligible. The model thus remains under-trained on long-range dependencies, even if in principle it has the capacity. This results in an effective context length much shorter than the maximum. Indeed, most open-source LLMs end up with effective context well under 50% of what they were trained for An et al. .
@@ 3.3 Positional Encoding Saturation and Long-Range Attenuation
Even if we had abundant training data for long contexts, the representation of position itself can become a limiting factor. Transformers rely on positional encodings to inject order information (since self-attention alone is order-invariant) Vaswani et al. (2017); Huang et al. (2023). However, standard positional encoding schemes have mathematical properties that hinder very long-range discrimination Touvron et al. (2023). This is called "positional encoding saturation", which means that beyond a certain context length, the model’s ability to distinguish positions or to maintain useful variability in positional signals deteriorates.
@@
Rotary Position Embedding (RoPE). Modern LLMs often use RoPE to encode relative positions by rotating query/key vectors. RoPE faces a similar issue as it encodes relative phase shifts such that the inner product of queryi and keyj implicitly includes a factor $\cos(\theta(i-j))$ (for each frequency band). Without adjustment, $\cos(\theta\Delta)$ becomes very small for large $\Delta$. This yields a “long-range attenuation” effect, and attention scores between tokens far apart are exponentially dampened. The base frequency used in RoPE effectively sets the length scale at which attention fades. If the base is not scaled for a larger window, beyond a certain distance, the overlap between rotated vectors is nearly zero. Empirical evidence of this comes from attempts to extend context windows: simply increasing a model’s maximum position with RoPE without rescaling leads to steep increases in perplexity and degraded utility beyond the original length Xiong et al. (2023); Bai et al. (2024a). Recent research identified this as a key limitation and proposed NTK-aware or scaled RoPE, which amplifies or adjusts the rotation frequency to “stretch” out the positional encoding over a longer span Xiong et al. (2023).
@@ 3.4 Attention Computation Limits: Softmax Dynamics and Memory Complexity
The third fundamental limitation lies in the attention mechanism’s computation itself when faced with extremely long sequences (as shown in Figure  5). The softmax self-attention in Transformers has a quadratic complexity in sequence length $N$ for both computation and memory, making very large $N$ costly Dao (2023). But beyond runtime, there are mathematical reasons why softmax attention struggles as $N$ grows, especially for reasoning tasks Li et al. (2023a); Liu et al. (2023a).
Softmax competition and diffusion of focus. In an attention head, the probability of attending to any given token is $\text{softmax}(e_{ij})=\frac{\exp(e_{ij})}{\sum_{k=1}^{N}\exp(e_{ik})}$, where $e_{ij}$ is the compatibility of token $i$’s query with token $j$’s key. As $N$ increases, the denominator grows with many terms Huang et al. (2023). If there is one relevant token among a sea of N1N-1 irrelevant ones, the model’s query must assign that relevant token a logit advantage on the order of $\log N$ to maintain a fixed attention probability. For example, to have 50% of the attention mass on one token out of $N$, the score for that token needs to be about $\ln(N)$ larger than the average of the others. This is a steep requirement: as context length grows, the model must sharpen the attention distribution more and more to pick out a single item. If the model’s scores for irrelevant tokens have some variance, a large $N$ increases the chance that some distractor token will get a moderately high score by coincidence, eating into the probability of the true relevant token. This effect can be viewed as a type of combinatorial noise which means that with many keys, the softmax normalization makes it difficult to preserve a strong signal for the correct one unless the model has learned extremely fine-grained, high-contrast scoring. In practice, as contexts lengthen, attention tends to diffuse, and it often spreads over many tokens or attends mostly to the recent segment, unless a very obvious keyword or cue is present to focus on the far context Liu et al. (2023c).
@@
Memory and precision. The quadratic memory use of attention means practical implementations struggle with very long inputs. Even if a model supports 100K tokens, performing attention on that many tokens can hit memory limits or require dumping to slower memory, which introduces numerical precision challenges. Additionally, summing over 100K exponentiated scores in softmax can lead to extremely large or small values, testing the limits of floating-point precision Dao (2023). Transformers process all tokens in parallel, which means every layer has to recompute interactions across the full sequence. With many layers, the opportunities for error accumulation or gradient diminishing over long ranges increase. In contrast, a recurrent process (like a state-space model) carries information forward iteratively, which has its own challenges (e.g., gradient vanishing through time) but uses a different mechanism for long-term dependency. In transformers, while skip connections help gradients propagate, there is no persistent memory that carries over from token to token beyond what attention redistributes at each layer. If at some intermediate layer the model fails to propagate a piece of information from position $j$ to $i$, later layers can only recover it if some indirect path exists. With very deep contexts, ensuring that all needed long-range links are formed somewhere in the stack is non-trivial.
Empirical benchmark results echo these theoretical concerns. $\infty$ Bench Zhang et al. (2024a); Chang et al. (2023); Pang et al. (2021); Bai et al. (2023); Kočiskỳ et al. (2018) which pushes context to 100K+ tokens, finds that current long-context LLMs “still require significant advancements” to handle 100K tokens effectively. In LongBench’s Bai et al. (2024b) multitask evaluation, models without explicit long-context training or architectural adjustments see dramatic drops in accuracy on tasks as input length increases, and as NeedleBench’s Li et al. (2025a) Ancestral Trace Challenge shows, even at a relatively modest 2K length, complex logical reasoning across dispersed information is often beyond the reach of today’s transformers. The limitations of attention become most apparent when naive long contexts meet tasks requiring synthesis of widely separated pieces; either the transformer tends to focus myopically on one part or gets lost trying to handle everything, revealing a fundamental lack of robust long-range reasoning.
@@
Left-skewed training over positions (“positional undertraining”) Model underuses late-context tokens; performance drops when evidence appears in the middle/end of long inputs. Data-side: Reweight or resample long-range pairs; curriculum that up-weights large relative distances; extend pretraining/fine-tuning on long sequences. Index remapping: Position redistribution/shift (e.g., StRing) to reuse well-trained low indices at inference. Loss shaping: depth-aware losses to boost gradients at large separations. Bai and others (2023); Li and others (2024); Wang and others (2024); Saito and others (2025)
@@
Positional encoding attenuation (sinusoidal & vanilla RoPE saturate) Dot products between distant positions vanish/oscillate; far tokens receive negligible positional reinforcement. RoPE scaling: NTK-aware or base-frequency scaled RoPE; dynamic/learned scaling per head. Relative schemes: ALiBi, T5-style relative attention, KERPLE; monotone distance biases that do not saturate; hybrid absolute+relative mixes. Content-aided: strengthen content cues via contrastive span objectives. Press et al. (2021); Su and others (2021); Peng and others (2023); Bai and others (2023)
@@
Softmax crowding (logit margin must grow like $\ln N$) Single relevant token drowned among $N$ distractors; attention mass diffuses as context grows. Architectural: Top-kk/sparse attention, landmark/sink tokens, hierarchical routing; multi-hop retrieval within the model. Training: margin-aware objectives on salient spans; hard-negative mining with long-context distractors; temperature scheduling. Inference: focused rereading (two-pass refine), constraint hints. Li and others (2024); Zhang and others (2024); Dao et al. (2022); Child et al. (2019)
@@
Quadratic memory/compute of global attention 128K–1M tokens infeasible or numerically brittle; KV cache and memory blow-ups. Algorithmic: FlashAttention/IO-aware kernels; block-sparse/dilated/linear-time variants; chunked sliding-window with cross-chunk summaries. System: paged KV cache, quantization of KV (NF4/FP8), CPU offloading with prefetch; windowed decoding. Dao et al. (2022); Child et al. (2019); Beltagy et al. (2020b); Dao (2023); Li et al. (2023a); Liu et al. (2023a)
@@
Fragmented long-range reasoning (missing multi-hop chains) Model fails when evidence is dispersed and must be combined over long spans. Planner–reader loops: iterative reasoning over retrieved segments; scratchpad/state passing across segments. Supervision: multi-hop chain objectives with distant supervision; compositional curricula. Structure: section headers/summaries injected as anchors. Bai and others (2023); Lewis et al. (2020b)
@@
Context selection is noisy (irrelevant filler dilutes signal) Performance collapses as irrelevant text increases density. Retrieval/compression: RAG; learned summarization/compression to evidence sketches; entropy/gradient-based token pruning; saliency pre-filters before LLM. Safety net: late fusion of multiple compressed candidates. Bai and others (2023); Li and others (2024); Rae et al. (2019); Lewis et al. (2020b)
@@
Index generalization beyond trained $L$ (learned embeddings) Positions beyond the seen range are near-initialization or ill-defined. Inter/extrapolation: continuous position functions; Fourier features with scale tuning; spline-based or low-rank position decoders; rope-based annealing during finetune. Su and others (2021); Peng and others (2023); Press et al. (2021)

### v2

@@ 2.3 Evaluation Misalignment and Guessing Incentives
The vast majority of present-day LLM evaluations employ binary grading. Responses are scored as either correct (1) or incorrect (0), with no partial credit for expressing uncertainty (Xu et al., 2024a). Under such schemes, abstentions, which can be responses like “I don’t know,” “I’m uncertain,” or “I cannot answer without more information”, receive the same zero score as confident but incorrect answers. This creates a rational incentive structure that strictly favors guessing over abstention.
@@
A meta-analysis of influential benchmarks confirms this pervasive issue (Yang et al., 2023). Among the most widely cited evaluations, namely MMLU-Pro, GPQA, Omni-MATH, IFEval, SWE-bench, MATH, BBH, HLE, all employ binary grading with no credit for uncertainty expressions (Li et al., 2023c). MMLU-Pro and GPQA are standard multiple-choice exams with no “I don’t know” option. Omni-MATH uses equivalence grading (often via LM judges) that compares outputs to ground-truth answers, assigning full credit for correctness and zero otherwise. IFEval grades instruction-following compliance programmatically, again binarized. Only WildBench, which evaluates real user chats on a 10-point scale, offers partial credit for uncertainty, but even there, the rubric suggests that “I don’t know” responses score lower than “fair” responses with minor hallucinations, thus still incentivizing guessing (Kirichenko et al., 2025).
@@
When models are trained via reinforcement learning from human feedback (RLHF) or direct preference optimization (DPO), the reward model inherits biases from human annotators who often prefer confident, fluent answers over hedged or uncertain ones (Xu et al., 2024a). Even when a model is unsure, producing a plausible-sounding fabrication receives higher reward than admitting ignorance. This creates a reward hacking dynamic, wherein models learn to maximize perceived confidence and fluency, which correlates with reward, rather than factual accuracy.
@@
where $\beta$ is the inverse temperature. If $R$ rewards verbosity and confidence over accuracy, the learned policy $\pi^{*}$ shifts toward hallucination-prone policies that generate detailed fabrications. This phenomenon is exacerbated by LM-as-judge evaluation, where another LLM grades outputs. LM judges are susceptible to length bias, favoring longer responses even when they contain errors, and often fail to detect subtle factual inaccuracies, grading incorrect but fluent hallucinations as correct (Yao et al., 2024a).
@@
In practice, LLMs exhibit significant overconfidence. They assign high probabilities to incorrect answers, i.e., $\mathbb{P}[\text{correct}\mid p_{\theta}(c)=0.9]\ll 0.9$. This miscalibration is particularly severe on long-tail queries and out-of-distribution inputs, where the model has minimal training signal (Sainz et al., 2023).
@@
where $\mathcal{R}_{\text{mem}}(h)\ll\mathcal{R}_{\text{true}}(h)$ reflects near-perfect recall of memorized answers. This distorts leaderboard rankings: models with higher contamination rates appear to hallucinate less, while their real-world performance remains poor (Xu et al., 2024a).
@@
Addressing evaluation misalignment requires major socio-technical reforms (Kirichenko et al., 2025). First, confidence-aware grading should assign partial credit for well-calibrated uncertainty expressions. Instead of binary $\{0,1\}$ scores, evaluations could use:
@@ 2.4 Creativity-Factuality Trade-off
LLMs generate text by sampling from a learned probability distribution $p_{\theta}(x_{t}\mid x_{<t})$ over next tokens. The sampling strategy governs the balance between exploitation (selecting high-probability, safe continuations) and exploration (venturing into lower-probability, generative territory). This trade-off is controlled by hyperparameters like temperature $T$ and nucleus (top-$p$) sampling (Peeperkorn et al., 2024).
@@
where $z_{t}$ is the logit for token $x_{t}$. As $T\to 0$, sampling becomes deterministic (greedy decoding), favoring the single most likely token. As $T\to\infty$, the distribution flattens toward uniform (high exploration, minimal exploitation). Empirically, low temperatures ($T\approx 0.1$-$0.5$) produce repetitive, conservative text with high factual accuracy but limited novelty. High temperatures ($T\approx 1.0$-$2.0$) yield diverse, creative outputs but also frequent hallucinations as the model samples unlikely tokens that lead to fabricated details (Nguyen et al., 2024).
@@
The challenge is distinguishing these cases. Models lack explicit mechanisms to estimate epistemic vs. aleatoric uncertainty, so sampling strategies cannot adapt. A fixed high temperature boosts creativity but also hallucination, while a fixed low temperature reduces hallucination but also stifles creativity (Farquhar et al., 2024).
@@
showing a negative correlation: improving creativity ($d\mathcal{C}>0$) necessitates reducing accuracy ($d\mathcal{A}<0$), and vice versa (Nguyen et al., 2024).
@@
A provocative theoretical perspective argues that hallucination is necessary for improvisation in LLMs (Jiang et al., 2024). If a model were constrained to produce only outputs directly supported by its training data, it could never generate truly novel text because every sequence then would be a recombination of observed patterns. Creativity demands extending beyond the training distribution, which by definition involves assigning non-zero probability to unseen or low-probability continuations. These exploratory samples constitute hallucinations when they introduce factual errors, but they are indispensable for creative generation. Formally, let $\mathcal{D}_{\text{train}}$ denote the support of the training distribution. A model that generates only from $\mathcal{D}_{\text{train}}$ produces zero hallucinations but also zero novel outputs. To improvise, the model must sample from regions where $p_{\text{train}}(x)\approx 0$ but $p_{\theta}(x)>0$. Whether such samples are creative or hallucinatory depends on the task: in fiction, they are celebrated as originality; in fact-based QA, they are condemned as fabrications. The line between creativity and hallucination blurs, with both stemming from the model’s learned ability to extrapolate beyond its training data.
@@ 3.1 Problem Description
Transformer-based LLMs have pushed context lengths into the tens or even hundreds of thousands of tokens, enabling applications like book summarization, multi-document QA, and long conversation memory (Vaswani et al., 2017; Beltagy et al., 2020a; Zaheer et al., 2020; OpenAI Achiam et al., 2023; Grattafiori et al., 2024; Liu et al., 2024a; Tomczak and Kuppannagari, 2025). Yet despite larger context windows, practical long-context reasoning often falls short of expectations (Qiu et al., 2020; Han et al., 2021). Models frequently cannot effectively utilize the full window. The effective usable context of many LLMs is much shorter than their nominal context length (Wang and others, 2024; Liu et al., 2023c; Wang and others, 2024; Huang et al., 2023). Even a 70B model trained to 128K context (Llama3.1) was found to only leverage about 64K effectively (Grattafiori et al., 2024; Wang and others, 2024) and in a comprehensive benchmark (LongBench) with average inputs  6.7K words, even a strong 16K-token model (GPT-3.5-Turbo-16k) still struggles on longer contexts (Radford et al., 2019; Brown et al., 2020; Wei et al., 2022). Clearly, simply extending the context window does not guarantee strong reasoning across that entire length. This review explores the reason why. We focus on three core factors (as shown in Figure 5), including training data positional distribution, positional encoding limits, and attention computation constraints, and explain how each fundamentally hinders long-range reasoning (Wang and others, 2024; Huang et al., 2023).
@@ 3.2 Left-Skewed Training Distribution and Undertraining of Long Positions
One root cause is the left-skewed distribution of token positions in training data (Wang and others, 2024). During pretraining, model inputs are far more likely to be shorter texts or early segments of long texts, rather than full-length sequences. This creates a severe imbalance: tokens at later positions (far right in the context) are extremely underrepresented (Xiong et al., 2023), as also shown in Figure 4. Recent analyses confirm this “left-skewed position frequency distribution” in large corpora (Wang and others, 2024; 2024). For example, in the SlimPajama dataset (a massive webtext corpus), the frequency of examples using very long-range token distances is vanishingly low: using a 2048 token window, less than 20% of the training pairs involve distances in the upper half of the window, and less than 5% involve the extreme end of the window (Bai et al., 2024a).
@@
In gradient-based training, rarely seen position interactions receive only tiny updates (Wang and others, 2024; Liu et al., 2023c). Intuitively, the model learns to predict the next token mainly using nearby context, and it gets much less practice using very distant context. Formally, one can consider the training loss as an expectation over position indices: if the probability of a position beyond, say, 75% of the max context is extremely small, then the contribution of those positions to the loss (and thus to gradient updates) is negligible. The model thus remains under-trained on long-range dependencies, even if in principle it has the capacity. This results in an effective context length much shorter than the maximum. Indeed, most open-source LLMs end up with effective context well under 50% of what they ostensibly trained for (Wang and others, 2024).
@@ 3.3 Positional Encoding Saturation and Long-Range Attenuation
Even if we had abundant training data for long contexts, the representation of position itself can become a limiting factor. Transformers rely on positional encodings to inject order information (since self-attention alone is order-invariant) (Vaswani et al., 2017; Huang et al., 2023). However, standard positional encoding schemes have mathematical properties that hinder very long-range discrimination (Touvron et al., 2023). This is called positional encoding saturation, which means that beyond a certain context length, the model’s ability to distinguish positions or to maintain useful variability in positional signals deteriorates.
@@
Rotary Position Embedding (RoPE). Modern LLMs often use RoPE to encode relative positions by rotating query/key vectors. RoPE faces a similar issue as it encodes relative phase shifts such that the inner product of queryi and keyj implicitly includes a factor $\cos(\theta(i-j))$ (for each frequency band). Without adjustment, $\cos(\theta\Delta)$ becomes very small for large $\Delta$. This yields a “long-range attenuation” effect, and attention scores between tokens far apart are exponentially dampened. The base frequency used in RoPE effectively sets the length scale at which attention fades. If the base is not scaled for a larger window, beyond a certain distance, the overlap between rotated vectors is nearly zero. Empirical evidence of this comes from attempts to extend context windows: simply increasing a model’s maximum position with RoPE without rescaling leads to steep increases in perplexity and degraded utility beyond the original length (Xiong et al., 2023; Bai et al., 2024a). Recent research identified this as a key limitation and proposed NTK-aware or scaled RoPE, which amplifies or adjusts the rotation frequency to “stretch” out the positional encoding over a longer span (Xiong et al., 2023).
@@ 3.4 Attention Computation Limits: Softmax Dynamics and Memory Complexity
The third fundamental limitation lies in the attention mechanism’s computation itself when faced with extremely long sequences (as shown in Figure  5). The softmax self-attention in Transformers has a quadratic complexity in sequence length $N$ for both computation and memory, making very large $N$ costly (Dao, 2023). But beyond runtime, there are mathematical reasons why softmax attention struggles as $N$ grows, especially for reasoning tasks (Li et al., 2023a; Liu et al., 2023a).
Softmax competition and diffusion of focus. In an attention head, the probability of attending to any given token is $\text{softmax}(e_{ij})=\frac{\exp(e_{ij})}{\sum_{k=1}^{N}\exp(e_{ik})}$, where $e_{ij}$ is the compatibility of token $i$’s query with token $j$’s key. As $N$ increases, the denominator grows with many terms (Huang et al., 2023). If there is one relevant token among a sea of N1N-1 irrelevant ones, the model’s query must assign that relevant token a logit advantage on the order of $\log N$ to maintain a fixed attention probability. For example, to have 50% of the attention mass on one token out of $N$, the score for that token needs to be about $\ln(N)$ larger than the average of the others. This is a steep requirement: as context length grows, the model must sharpen the attention distribution more and more to pick out a single item. If the model’s scores for irrelevant tokens have some variance, a large $N$ increases the chance that some distractor token will get a moderately high score by coincidence, eating into the probability of the true relevant token. This effect can be viewed as a type of combinatorial noise which means that with many keys, the softmax normalization makes it difficult to preserve a strong signal for the correct one unless the model has learned extremely fine-grained, high-contrast scoring. In practice, as contexts lengthen, attention tends to diffuse, and it often spreads over many tokens or attends mostly to the recent segment, unless a very obvious keyword or cue is present to focus on the far context (Liu et al., 2023c).
@@
Memory and precision. The quadratic memory use of attention means practical implementations struggle with very long inputs. Even if a model supports 100K tokens, performing attention on that many tokens can hit memory limits or require dumping to slower memory, which introduces numerical precision challenges. Additionally, summing over 100K exponentiated scores in softmax can lead to extremely large or small values, testing the limits of floating-point precision (Dao, 2023). Transformers process all tokens in parallel, which means every layer has to recompute interactions across the full sequence. With many layers, the opportunities for error accumulation or gradient diminishing over long ranges increase. In contrast, a recurrent process (like a state-space model) carries information forward iteratively, which has its own challenges (e.g., gradient vanishing through time) but uses a different mechanism for long-term dependency. In transformers, while skip connections help gradients propagate, there is no persistent memory that carries over from token to token beyond what attention redistributes at each layer. If at some intermediate layer the model fails to propagate a piece of information from position $j$ to $i$, later layers can only recover it if some indirect path exists. With very deep contexts, ensuring that all needed long-range links are formed somewhere in the stack is non-trivial.
Empirical benchmark results echo these theoretical concerns. $\infty$ Bench (Zhang et al., 2024a; Chang et al., 2023; Pang et al., 2021; Bai et al., 2024b; Kočiskỳ et al., 2018) which pushes context to 100K+ tokens, finds that current long-context LLMs “still require significant advancements” to handle 100K tokens effectively. In LongBench’s (Bai et al., 2024b) multitask evaluation, models without explicit long-context training or architectural adjustments see dramatic drops in accuracy on tasks as input length increases, and as NeedleBench’s (Li and others, 2024) Ancestral Trace Challenge shows, even at a relatively modest 2K length, complex logical reasoning across dispersed information is often beyond the reach of today’s transformers. The limitations of attention become most apparent when naive long contexts meet tasks requiring synthesis of widely separated pieces; either the transformer tends to focus myopically on one part or gets lost trying to handle everything, revealing a fundamental lack of robust long-range reasoning.
@@
Left-skewed training over positions (“positional undertraining”) Model underuses late-context tokens; performance drops when evidence appears in the middle/end of long inputs. Data-side: Reweight or resample long-range pairs; curriculum that up-weights large relative distances; extend pretraining/fine-tuning on long sequences. Index remapping: Position redistribution/shift (e.g., StRing) to reuse well-trained low indices at inference. Loss shaping: depth-aware losses to boost gradients at large separations. (Bai et al., 2024b; Li and others, 2024; Wang and others, 2024; Saito and others, 2025)
@@
Positional encoding attenuation (sinusoidal & vanilla RoPE saturate) Dot products between distant positions vanish/oscillate; far tokens receive negligible positional reinforcement. RoPE scaling: NTK-aware or base-frequency scaled RoPE; dynamic/learned scaling per head. Relative schemes: ALiBi, T5-style relative attention, KERPLE; monotone distance biases that do not saturate; hybrid absolute+relative mixes. Content-aided: strengthen content cues via contrastive span objectives. (Press et al., 2021; Su and others, 2021; Peng and others, 2023; Bai et al., 2024b)
@@
Softmax crowding (logit margin must grow like $\ln N$) Single relevant token drowned among $N$ distractors; attention mass diffuses as context grows. Architectural: Top-kk/sparse attention, landmark/sink tokens, hierarchical routing; multi-hop retrieval within the model. Training: margin-aware objectives on salient spans; hard-negative mining with long-context distractors; temperature scheduling. Inference: focused rereading (two-pass refine), constraint hints. (Li and others, 2024; Zhang and others, 2024; Dao et al., 2022; Child et al., 2019)
@@
Quadratic memory/compute of global attention 128K–1M tokens infeasible or numerically brittle; KV cache and memory blow-ups. Algorithmic: FlashAttention/IO-aware kernels; block-sparse/dilated/linear-time variants; chunked sliding-window with cross-chunk summaries. System: paged KV cache, quantization of KV (NF4/FP8), CPU offloading with prefetch; windowed decoding. (Dao et al., 2022; Child et al., 2019; Beltagy et al., 2020b; Dao, 2023; Li et al., 2023a; Liu et al., 2023a)
@@
Fragmented long-range reasoning (missing multi-hop chains) Model fails when evidence is dispersed and must be combined over long spans. Planner–reader loops: iterative reasoning over retrieved segments; scratchpad/state passing across segments. Supervision: multi-hop chain objectives with distant supervision; compositional curricula. Structure: section headers/summaries injected as anchors. (Bai et al., 2024b; Lewis et al., 2020b)
@@
Context selection is noisy (irrelevant filler dilutes signal) Performance collapses as irrelevant text increases density. Retrieval/compression: RAG; learned summarization/compression to evidence sketches; entropy/gradient-based token pruning; saliency pre-filters before LLM. Safety net: late fusion of multiple compressed candidates. (Bai et al., 2024b; Li and others, 2024; Rae et al., 2019; Lewis et al., 2020b)
@@
Index generalization beyond trained $L$ (learned embeddings) Positions beyond the seen range are near-initialization or ill-defined. Inter/extrapolation: continuous position functions; Fourier features with scale tuning; spline-based or low-rank position decoders; rope-based annealing during finetune. (Su and others, 2021; Peng and others, 2023; Press et al., 2021)